How many mililiters of 0.75% sugar solution must be added to 100Ml of 1.5% sugar solution to form a 1.25%?
Answer by milo1954(10)

0.75%X + 1.5% 100 = (100+X)1.25%
0.75X+1.5*100=1.25(X+100)
0.75X+150=1.25X+125
1.25X+125=0.75X+150
1.25X-1.75X=150-125
0.5X=25
X=50